Please help me find a solution to this problem I am having trouble understanding how to work it out. If the volume of a sphere is 500Pi/3m cubed what is the diameter of the sphere?
Answer by scott8148(6628)

volume of a sphere=(4/3)(pi)(r^3)

(4/3)(pi)(r^3)=500(pi)/3 __ dividing by (4/3)pi __ r^3=125

taking cube root __ r=5

d=2r __ d=10